Grade 11General Physics

The length of seconds pendulum at a place is 1.02m. Find the time period of another pendulum of 0.51m length at the same place.

Pavani Jaladi
8 Years agoGrade 11

1 Answer

Athul Joseph
8 Years ago
T=2*pi*(l/g)^0.5.Here T is directly proportional to (l)^0.5.T/t=(1.02÷0.51)^0.5t=T/(2^0.5).where T is the time period when the length is 1.02 and t is the time period where length is 0.51.